OpenStudy (anonymous):

what is the answer to five and two thirds minus one and 1 fifths

OpenStudy (fanduekisses):

4 and 7/15 ?

OpenStudy (anonymous):

\[5\frac{ 2 }{ 3 } - 1\frac{ 1 }{ 5 }\] You'd make sure that they have the same denominator first. This means multiplying the 2/3 by five on the bottom and the top, and multiplying the 1/5 by 3 on the bottom and the top. \[5\frac{ 10 }{ 15 }-1\frac{ 3 }{ 15 }\] First you would subtract the fractions, then the numbers before them. This would give you \[4\frac{ 7 }{ 15 }\] There is another way. That just deals with only the fraction and not the whole number. If you are interested in seeing that as well then let me know.